OpenStudy (anonymous):

-x - y + (x^2 + y^2)/(x-y) simplify

OpenStudy (anonymous):

try clearing the fraction by multiplying everything by (x-y)

OpenStudy (anonymous):

-x - y + ((x^2 + y^2)/(x-y) ) I rewrite is this righ?

OpenStudy (anonymous):

(2y^2)/x-y

OpenStudy (anonymous):

yes it is nancy

OpenStudy (anonymous):

end up with (-x^2-y)-(x-y^2)+(x^2+y^2)

OpenStudy (anonymous):

LCD: x-y

OpenStudy (anonymous):

(-x-y)+(x^2+y^2)/(x-y) =[(-x-y)(x-y)/(x-y)]+[(x^2+y^2)/(x-y)]

OpenStudy (anonymous):

Maph wouldn't the point of using the LCD be to clear the fraction?

OpenStudy (anonymous):

time:(x-y)(-x - y)=-x^2 -xy +xy +y^2 simplyfy = -x^2 + y^2 (-x^2+y^2+x^2+y^2)/(x-y) simplyfy (2y^2)/(x-y)